If 3.56g of butane is burned underneath a cup holding 1.00L of water at 21.0
1
Expert's answer
2021-08-23T02:56:25-0400

2C4H10 + 13O2 → 8CO2 + 10H2O


Moles of butane "=\frac{3.56}
{58.12}=0.06moles"


Moles of water"=\frac{10}{2}×0.06=0.3moles"


Molarity of water"=\frac{0.3}{1.00}=0.3M"